# Hard Contacts with Soft Gradients: Refining Differentiable Simulators for Learning and Control

### v1

@@ 3.2 Adaptive stepsize integration: Numerical precision on demand
A standard methods for integrating ODEs with variable stiffness are adaptive integrators. The idea behind adaptive stepsize integration is simple: Two numerical integrators of different order compute the next state. Their difference provides an estimate of the error. If the error is smaller than a given threshold, the step is accepted, otherwise the step is rejected and the procedure is repeated with a different stepsize chosen by a feedback controller. For further details on the rich history of adaptive stepsize integration, see e.g. [21, 22, 23, 24]. We use Diffrax [25] for efficient numerical integration in JAX, taking advantage of its solver flexibility and enhanced backpropagation modes. We extend it to integrate quaternions and stateful actuators seamlessly and make Diffrax compatible with MJX. Further details are provided in Suppl. B.2.
@@
Using an adaptive integrator with MJX eliminates oscillations in the bounce example. However, gradients for some object primitives (capsule, cylinder, box) experience offsets due to non-differentiable operations in the collision detector arising from discrete case distinctions. We smoothened them with standard proxies, leading to results in the bottom row of Fig. 5, where analytical gradients nearly match central differences. Henceforth, we refer to MJX with the Diffrax integrator and smoothened collision detection as DiffMJX.
@@
Figure 8: Billiard simulation. Top: Force $F$ acts on the white ball to minimize the loss.
Bottom: Despite the loss derivative being zero if the balls do not collide, DiffMJX with CFD provides informative gradient information.
4 Contact force gradients from a distance
We can now compute the correct gradients of the dynamics. However, are these gradients also informative? For example, consider a simulated billiard table setup as shown in Fig. 8. The white ball should be shoot to hit the black ball and make it reach the target. We are exerting a force $F$ onto the white ball in the first timestep and would like to optimize this force to minimize the distance $L$ between the black ball and the target. As long as the balls collide, MJX with adaptive integration yields informative gradients $\nabla_{F}L$. However, if $F$ does not cause the balls to touch, then $\nabla_{F}L$ is zero and therefore uninformative for optimization. Therefore, we propose contacts from distance (CFD), a method for computing contact forces for positive signed distances $r$ in penalty-based simulation to yield informative gradients even if objects are not in contact.
Creating artificial contact forces from a distance.
As discussed in Suppl. 2, the magnitude of contact forces in MuJoCo is determined by the impedance $d(r)$ and position-level reference acceleration $h(r)$. To enable the solver to apply CFD, we can augment $d(r)$ as depicted in Fig. 8. Here, we do not alter $d(r)$ for $r<0$, but instead extend $d(r)$ for $r>0$. This continuation is parametrized by solimp-CFD parameters $(d_{c},d_{0},w_{c},m_{c},p_{c})$. By default, the curve smoothly continues MuJoCo’s impedance at $d_{0}$ and tapers of to $d_{c}=0$ to ensure smooth differentiability. The CFD width $w_{c}$ specifies the distance for which artificial contact forces are generated and, in our experiments, has been varied between 1 cm and 1 m. Moreover, we soften the reference acceleration $h(r)$ by replacing the ReLU function on the signed distance with a softplus (Fig. 8). This yields modified contact forces $f_{\text{CFD}}$, and hence the modified ODE $\dot{v}_{\text{CFD}}=M^{-1}\left(\tau-c+J^{\top}f_{\text{CFD}}\right)$.
@@ 4 Contact force gradients from a distance
CFD provides contact gradients for positive signed distances at the expense of physical realism. As shown in Fig. 8 (top), naive simulation with CFD results in floating objects. In this example, CFD can be seen as introducing a compliant layer mimicking a soft foam mat of thickness $w_{c}$ being placed atop the actual surface. In turn, the quadrupedal robot appears to hover above the surface. As significantly altering simulation realism is not an option, we are faced with the question: Can CFD be used to obtain informative contact gradients without affecting simulation realism?
To positively answer this question, we propose a simple idea: we keep the forward computation (vanilla MJX or DiffMJX), but enable CFD only in a simulation used for gradient computation via automatic differentiation. This is achieved by a variation of the straight-through-trick applied to the forward dynamics model, as detailed in Suppl. B.3. The resulting procedure is illustrated in Fig. 8 (bottom). Note that this approach requires the simulator’s forward pass to be computed twice. However, the gradient, which typically dominates the computational cost, is still only evaluated once.
Revisiting the billiard example from the beginning of this section (Fig. 8), we see that by using the straight-through-trick, DiffMJX computes both the correct loss as well as gradients pointing towards the loss minimum despite the balls not being in contact.
Figure 9: Left: Estimation of a cube’s side length in MJX via gradient descent using multi-step ahead predictions. Right: Experimental setup for collecting cube toss data. Image adapted from [26].
5 Evaluation
5.1 System Identification
Parameter identification in the presence of hard contacts remains a laborious task. If contacts are hard, even learning the dynamics of a cube requires impractical amounts of data for “naive” neural network regression [27]. In comparison, penalty-based simulators can capture hard contacts, but the lack of correct gradients hinders efficient parameter estimation [28]. Therefore, recent work introduced intricate analytical pipelines for cube geometry estimation [26, 29] and graph-based networks for learning contact dynamics [30]. In what follows, we use the same real-world data as used in [26, 29, 30]. We demonstrate that DiffMJX with CDF enables simulator parameter estimation via standard gradient-based optimization.
Dataset and training setup.
We use the Contactnets dataset [26] which consists of $550$ trajectories of an 10 cm acrylic cube that has been repeatedly tossed onto a wooden table. For training, trajectories are split into segments of length five such that the simulator is tasked to unroll four future steps starting from the initial state. Each segment and its prediction is fed to an $L_{2}$ loss whose gradient is used for gradient-based optimization using Adam [31]. For systems with stiff dynamics, we favor multi-step-ahead predictions over one-step-ahead predictions, as they capture the cumulative effects of prediction errors over time. This setup enables a fairer analysis of MJX without CFD, as even for too small side length estimates, future state predictions can make contact to inform the optimization.
Training results.
The training results are shown in Fig. 9. Surprisingly, the refined version of MJX and MJX with CFD already achieve good estimation results with an error of around 5% relative to the ground-truth. If the side length is initialized at 60 mm or 140 mm, training either stalls fully or convergence is severely limited for MJX. The incorporation of CFD into MJX addresses convergence issues arising from poor initial parameters, while the integration of adaptive integration via DiffMJX significantly enhances estimation accuracy. DiffMJX improves estimation accuracy by dynamically adjusting the time steps during collisions, thereby mitigating time discretization errors. Further details and parameter estimation experiments are provided in Suppl. C.2. To the best of our knowledge, we are the first to demonstrate parameter estimation of real-world cube dynamics using an automatically differentiable penalty-based simulator. While this represents a promising step forward, further experimentation is necessary to fully characterize the scope and limitations of this approach.
5.2 Model Predictive Control
Next, we conduct experiments on gradient-based model-predictive control. We use a simple MPC loop in which, at every plan step, we refine a sequence of controls over a $256$-step horizon. In the gradient-based planner, we compute gradients by backpropagating the differentiable cost computed on the rollout of the current plan through the MJX simulator. The plan is then iteratively optimized using the Adam optimizer with a learning rate of $0.01$ for $32$ iterations. Finally, the resulting plan is executed for $16$ steps in simulation, after which the planning procedure is repeated with the previous plan as a warm start. As a simple baseline, we include a version of the predictive sampling planner from Mujoco MPC [32], which at every plan step samples $k=\{64,256,1024\}$ trajectories, and executes the lowest-cost plan. We significantly improved the performance of this planner for the muscular systems by resorting to brown noise [33] for sampling.
Models.
All our MPC experiments revolve around the muscle-tendon models provided by MyoSuite [34, 35]. Models include the MyoHand modified from the MyoChallenge 2022, which is comprised of 29 bones, 23 joints, and 39 muscle-tendon units. We also use a bionic model modified from the MyoChallenge 2024, which is comprised of the MyoArm with 27 degrees of freedom and 63 muscle-tendon units, and the simulated modular prosthetic limb with 26 degrees of freedom and 17 motor control units.
Figure 10: Left: Simulation cost evolution of gradient-based MPC, with and without contacts from distance (CFD), vs sampling-based MPC. The number for sampling indicates the number of samples used per planning step. Sampling has difficulties solving the dexterous in-hand manipulation task; gradients without CFD cannot solve the bionic tennis task. Right: Rendering of sampling-based MPC (1024 samples) vs gradient-based MPC with CFD on in-hand manipulation task. The goal is to swap the balls; the MyoHand model is actuated by 39 muscle-tendon units.
Dexterous in-hand manipulation: Determining crucial components in the MPC loop.
First, we consider an in-hand manipulation task, where the goal is to swap two balls in the MyoHand. The cost is given by the Euclidean distance between each of the balls and the respective target location, fixed in the frame of the hand.
The results are reported in Fig. 10. We find that gradient-based MPC can reliably solve this task, in contrast to the sampling-based planner. Overparameterization in the muscle-tendon model with at least two muscles per joint benefits the gradient-based planner by helping escape local minima, similar to its role in optimizing overparametrized neural networks. In contrast, RL and sampling-based planners struggle with scaling in overparametrized higher-dimensional systems [36]. First-order methods using differentiable simulation should be able to tackle more complex control problems.
Notably, this task does not require distant contacts because hand-ball interactions are frequent due to gravity. Moreover, we identify two crucial components of the gradient-based MPC loop: First is gradient clipping, which is important as the scale of gradients changes massively in the presence of contacts. This technique has also been reported to be effective in previous works on differentiable simulation [37, 38]. Second, we store the rollout cost of all gradient iterations and select the one with minimal cost. This is important as the cost landscape is highly non-convex, which is reflected in the non-monotonic cost evolution between the iterations of a planning step.
Finally, we test a more complex custom bionic tennis task on the bionic model. The task is to move a ball that is initially moving sideways to a target location below. This can be achieved by bouncing it back using a racket that is statically welded to the prosthetic hand, and then catching it at the target location with the muscle hand. In this task, the only cost supervision is the Euclidean distance of the ball to the target, the complicated sequential movement has to be discovered purely from this signal.
We report our findings in Fig. 10, see Fig. 1 for a rendering. By design, the task initialization is such that the ball misses both hands, hence this task is not solvable by purely gradient-based MPC using vanilla MJX. On the other hand, we observe that adding the contacts from the distance mechanism allows solving this task. The sampling-based planner is a strong baseline in this task and gets very close to solving it. Initially bouncing the ball back to the target only requires controlling the prosthetic arm which is relatively low-dimensional, hence the sampling-based planner achieves this part easily. However, as seen in the in-hand manipulation task, it struggles with precise control of the high-dimensional MyoHand.
6 Conclusion
In this work, we tackle the challenge of inaccurate gradient computation in differentiable penalty-based simulators, such as MuJoCo XLA (MJX), arising from hard contact dynamics. Through a 1D toy example, we demonstrate that these errors are a consequence of time-discretization and can be mitigated by reducing the integration stepsize. While reducing the stepsize improves gradient accuracy, it comes at the cost of increased simulation time and GPU memory usage. To address these limitations, we introduced DiffMJX, which is a combination of MJX with the adaptive numerical integration library Diffrax. This approach enables the use of adaptive time integration to reduce discretization errors while significantly reducing memory overhead through checkpointing. To further improve differentiability in contact-rich environments, we smoothen MJX’s collision detection pipeline, enabling accurate gradient computation of MuJoCo’s basic shape primitives for enabling stiff contact settings. Complementing DiffMJX, we propose Contacts from Distance (CFD), a technique that improves gradient utility by introducing small virtual contact forces between near-interacting bodies, without affecting the forward simulation. CFD is implemented via the straight-through-trick. We validate these contributions in a real-world system identification experiment, where DiffMJX successfully estimates the geometric parameters of a cube via simple gradient descent. Moreover, this work is the first to demonstrate that automatic differentiation with CFD can outmatch sample-based predictive planning in bionic dexterous manipulation tasks.
7 Limitations
7.1 Computational cost of DiffMJX
DiffMJX increases the computational cost of MJX depending on the integrator settings. A detailed analysis of DiffMJX computational cost is given in Figure S6. If an adaptive integrator is chosen, then the computational cost may notably increase as many more steps are taken during simulation. However, as we show that achieving well-behaved gradients in the presence of hard contacts requires a sufficiently small stepsize, this is a necessary price to pay. The alternative is to reduce the native fixed MJX stepsize, which will require many more steps when the stiffness of the ODE changes a lot due to collision events. Notably, there is a non-negligible increase in JIT compilation time when using Diffrax, as the program is compiled for the maximum number of ODE steps that is set by the user. That said, thanks to Diffrax’s advanced techniques for checkpointing, DiffMJX memory consumption is significantly reduced compared to MJX.
Through Diffrax, DiffMJX provides a whole suite of additional tuning knobs that can be used to improve gradient computation. The most important parameter by far is the integrator’s error tolerances, which directly determine to what extent discretization errors cause oscillations in the gradient signal. Ideally, we want to choose the smallest error tolerance that provides sufficiently good gradients to keep the computational footprint small. However, this error tolerance depends on the stiffness of the simulation, aka the hardness of contacts and joint limits. The error tolerance is an additional hyperparameter that requires tuning.
7.2 Impact of number of contacts on CFD
The documentation of MuJoCo XLA [^1] https://mujoco.readthedocs.io/en/stable/mjx.html#mjx-the-sharp-bits advises to use meshes with 200 vertices or fewer due to computational limitations on the current implementation of collision detection. If MJX is used with CFD, then every contact is added to the contact solver for which $r<w_{c}$. In turn, for dense meshes and large $w_{c}$, computation can become prohibitively large. The parameter $w_{c}$ can be made arbitrarily small such that, in its limit for $w_{c}=0$, we retain vanilla MJX with our improvements on the differentiability of the collision detector.
While gradients computed with CFD enable controller synthesis for highly-dimensional systems such as muscular robots, it can only do so effectively if the number of contacts remains small. That said, for in-hand manipulation tasks and robot locomotion tasks, the number of objects colliding at one instance is usually small enough to render CFD a useful tool for easing optimization with penalty-based simulators.
7.3 System Identification
For the very large initial cube side length initial value of $140$ mm, DiffMJX saw a significant drop in computation speed forcing us to abort these runs. This is not surprising as penetrations amounting to 40% of the total cube’s width cause large contact forces that significantly stiffen the ODE. In Appendix C.2, we reduce the integration error tolerance which allowed optimization to converge even for large initial parameters. Alternatively, one could speed up training for these too large geometry parameter initializations by first using a soft impedance that during training becomes annealed to become increasingly stiff.
In principle, our implementation of DiffMJX enables gradient computation with respect to any differentiable simulator parameter, including the cube’s mass and inertia, as well as contact model parameters such as solref and solimp. However, since gradient descent is a local optimization method, it is susceptible to convergence to local minima as shown in Appendix C.2. The results shown in Fig. 9 converged reliably when the simulated cube’s mass was set to the ground-truth value and the contact parameters were tuned to somewhat resemble the real-world motion. In scenarios where no prior knowledge of simulator parameters is available, one likely needs to combine DiffMJX with global optimization methods such as randomly sampling initial simulator parameters.
7.4 Gradient-based vs sampling-based MPC
In general, gradient computation with MJX via autodifferentiation is considerably slower than running only the forward simulation. This makes zeroth-order methods such as predictive sampling favorable whenever the task can be solved with them, as the runtime is much lower. However, sampling suffers from the curse of dimensionality, whereas gradients scale to extremely high-dimensional control problems.
For MPC, our CFD mechanism tends to be effective when the object needs to be pushed in a specific direction by the hand, but it performs poorly in grasping tasks. This limitation arises because, at a distance, the cumulative force arising from the hand’s CFDs conglomerate into a force that pushes objects away. However, the gradient does not encode the possibility of grasping, as this signal only emerges when the object is inside the hand. Fundamentally, this issue reflects the non-convex nature of the optimization landscape.
However, we do not believe that this is a problem that should be solved using a CFD mechanism; rather, in such tasks, the best approach would be to combine CFD with sampling to overcome the non-convexity while maintaining the favorable scaling of gradients with dimensionality.
Acknowledgements
The authors thank Onur Beker, Thomas Rupf, and Nico Gürtler for providing valuable feedback.
Supplementary Material for:
Hard Contacts with Soft Gradients: Refining Differentiable Simulators for Learning and Control
Appendix A Related Work
Differentiable simulators are an active area of research spanning multiple fields of physics, including elastic object and fluids modeling [40, 39, 7, 41] and modeling rigid body collisions [6, 7, 42, 18] (see [43] for a recent overview). As outlined in Figure S1, we focus on robotics simulators involving hard contact collisions, with the hierarchical objectives of (i) accurately simulating dynamics, (ii) computing correct simulator gradients, and (iii) obtaining informative gradients between non-colliding objects.
Figure S1: Overview on hierarchical goals that need to be accomplished to use simulator gradients for robot controller synthesis.
Simulating contact dynamics.
Commonly deployed robot simulators can be categorized according to their contact model into impulse-based [44, 45, 7], complementary-based [42, 46], and penalty-based [6] approaches. Due to their simplicity and speed, impulse-based simulators [47], such as those using DiffTaichi [48, 7], are widely used in game development but remain currently uncommon in robot controller synthesis. Complementary-based contact models compute constraint forces as a solution to a constrained optimization problem either in the form of a nonlinear-complementary problem (NCP) [42] or linear-complementary problem (LCP) [16, 15, 49, 50]. Exactly solving NCP is an NP-hard problem [51] that is considerably difficult to solve even approximately. The NCP formulation in Dojo is physically more accurate than LCP formulations [42], but lacks support for parallel computation limiting its utility for controller synthesis. That said, it is currently not clear which complementary formulation is best suited for robotics.
To improve computational efficiency, MuJoCo [6, 12] reformulates the complementary constraint problem as a convex optimization problem. This reformulation builds on prior work on complementarity-free approaches [52, 53]. A detailed description of MuJoCo is provided in Sec. B.1. Briefly, contact forces are computed by solving a global optimization problem, in which the ability of a constraint to generate force is modulated by the geometric penetration depth at the contact interface. As an alternative, Drake [16] also adopts a soft patch contact model [54, 55]. In turn, these approaches to contact resolution soften the dynamics.
Differentiating through contacts.
The choice of contact model has significant implications on the differentiability of the simulator. If complementary-based solvers exactly compute hard contact forces, then the resulting jumps in the dynamics aggravate gradient computation. Hence, recent literature on differentiating through complementary-based solvers propose analytical reformulations of the dynamics, e.g. based on the implicit function theorem [46, 42], or resort to randomized smoothing [56, 57, 58, 59, 60]. While differing in their computational cost, these approaches to smoothing have been shown to be empirically equivalent [55, 8].
While the dynamics of penalty-based simulators are inherently smooth, gradient inaccuracies introduced by time discretization grow with both contact stiffness and integration step size. Prior work has attributed this problem to time-of-impact discretization errors [7, 8]. As a result, accurately simulating near-rigid contacts within penalty-based frameworks requires significantly smaller integration step sizes, which limits their practicality for robot controller synthesis. In this work, we propose an alternative direction: adaptively integrating compliant simulators.
Contact-invariant optimization.
To guide an optimizer toward using body collisions for task facilitation, recent works in controller synthesis and path planning incorporate inter-body distances into the optimization objective. In RL, [61] added inter-point distances to rewards using a softmax function. Alternatively, the algorithms in MuJoCo MPC [32] rely on inter-point distances combined with trajectory samples establishing contact to inform the optimization. However, depending on the task at hand, setting up distance-based loss terms to inform an optimizer about the necessity for specific object collisions can quickly become cumbersome. As an alternative approach, Mordatch et al. [10] proposed the framework of “contact-invariant optimization” (CIO) in which a simulation is adjusted to also apply contact forces between non-colliding bodies. While the simulation can apply non-physical contact forces, the optimization is encouraged through the addition of several loss terms to minimize the usage of these forces. In [9], CIO was extended to in-hand manipulation of objects. Our proposed extension of contacts from distance (CFD) is inspired by CIO. Yet, while CIO applies artificial contact forces in the forward simulation, CFD resorts to the straight-through-trick to only exert contact forces in a simulation that is solely used for gradient computation. In turn, CFD does not require the addition of regularization terms to a loss when used for planning. Recently, [62] proposed soft signed distance fields as an alternative geometry representation, while also adjusting collision detection and contact force computation to be inherently soft. This approach which routes at the geometry level of robot simulation can be seen as an alternative avenue towards CIO.
Straight-through estimation.
The use of the straight-through-trick in CFD is inspired by a plethora of works in robotics and beyond. The code underlying the straight-through-trick is reported in JAX [63] documentation under the collective heading of straight-through estimation. Straight-through estimation has been used in [64], where the backward pass of a Heaviside function is replaced with an identity function to obtain sparse RNNs for model-based RL. Horuz et al. [75] demonstrate that using a ReLu activation with a custom backward function rivals smooth ReLu surrogates. Recently, [65] replaced simulator dynamics obtained from IsaacGym [17] with single-rigid body dynamics for the gradient computation in the backward pass. This approach is closely related to our work, but unlike CFD, [17] does not utilize the forward pass simulation in the backward pass, limiting it’s utility for complex control synthesis tasks and system identification.
Appendix B Method Details
B.1 Contact force computation in MuJoCo
A detailed description of MuJoCo is given in its documentation [^2] https://mujoco.readthedocs.io/en/stable/computation/index.html, here we merely provide a digestible summary of the key equations. Through the Newton-Euler equations of the unconstrained system (1), one can determine the acceleration produced by a specific generalized force, and utilizing the collision detector along with geometrical data, we can identify the location of contact forces exerted by body points. This leaves us with the question on the direction and magnitude of constraint forces. MuJoCo finds an approximate answer to this question by resorting to Gauss principle [19, 20] which states that nature chooses the acceleration of rigid-body systems such that the weighted square norm $\|J^{\top}f\|_{M}^{2}$ is minimized while all constraints are met exactly. MuJoCo deploys softened versions of Gauss principle and Baumgarte stabilization, reading
$\begin{split}(\dot{v},\dot{\omega})&=\underbrace{\underset{(x,y)}{\arg\min}\bigl{\|}x-M^{-1}(\tau-c)\bigr{\|}_{M}^{2}}_{\text{Inspired by Gauss' principle}}\quad+\underbrace{\bigl{\|}y-a_{\mathrm{ref}}\underset{}{\bigr{\|}_{R^{-1}}^{\text{Huber}}}}_{\text{Inspired by Baumgarte stabilization}}\\ &\qquad\text{subject to}\quad J_{\mathcal{E}}x-y_{e}=0,\quad J_{\mathcal{F}}x-y_{\mathcal{F}}=0,\quad J_{\mathcal{C}}\,x-y_{\mathcal{C}}\in\mathcal{K}^{*}\end{split}$ (S1)
with the index sets for equality, friction, and contact constraints $\mathcal{E},\mathcal{F},\mathcal{C}$, the friction cone (dual) $\mathcal{K}^{*}$, the regularizer $R>0$, and $\|\cdot\|^{\text{Huber}}$ denoting the Huber norm. The reference acceleration $a_{\mathrm{ref}}$ denotes the solver’s target for the constraint space acceleration $\dot{\omega}$. Drawing inspiration from Baumgarte stabilization [66], MuJoCo chooses $a_{\mathrm{ref}}$ as the dynamics of a spring-damper system, writing
$a_{\mathrm{ref},i}=-b_{i}(Jv)_{i}-k_{i}r_{i}=-\frac{2}{(d_{w}\cdot t_{c})}(Jv)_{i}-\frac{d(r_{i})}{(d_{w}^{2}\cdot t_{c}^{2}\cdot\phi_{d}^{2})}r_{i},$ (S2)
whose dynamics are determined by the impedance $d(r)$, and the solref parameters consisting of the time constant $t_{c}$ and the damping ratio $\phi_{d}$. The impedance $d(r)$ is the central function for determining constraint forces (such as contact forces). As illustrated in Figure 3, the impedance $d\in(0,1)$ is a function of the constraint violation $r$ and is specified by the solimp parameters $(d_{o},d_{w},w,\text{midpoint},\text{power})$ which determine its shape as a polynomial spline. MuJoCo’s documentation refers to the impedance as a “constraints ability to generate force” as it not only determines $a_{\mathrm{ref}}$, but also directly weights the cost for applying constraint forces in Eq. S1 via $R$, which is computed as
$R=\frac{1-d_{i}}{d_{i}}\hat{A}_{ii},$ (S3)
where $\hat{A}_{ii}$ denotes an approximation to the diagonal entries of the inverse inertia in constraint space $A_{ii}=JM^{-1}J^{\top}$. In turn, the constraint is hard if $R\rightarrow 0$, and more soft for larger $R$, approaching an infinitely soft (i.e. non-existent) constraint in the limit $R\rightarrow\infty$. As outlined in Appendix B.1, MuJoCo transforms Eq. S1 into a convex problem that is solved efficiently by an exact Newton method.
If $\mathcal{K}^{*}$ solely contains pyramidal or elliptic cone constraints, then the constrained optimization problem can be reduced to the convex problem
[…]
where the analytical function $s(\cdot)$ acts as a soft constraint penalty. By default, MuJoCo solves (S4) using an exact Newton method. While we do not delve into the inner workings of MuJoCo’s constraint solver, the reader should keep in mind that the number of solver iterations and its error tolerance affect the stability of the simulation.
B.2 ODE differentiation in Diffrax
Diffrax: Numerical integration with JAX
The ODE solved by the Mujoco simulator can be written as
$x(0)$ $=x_{0}$ $\dot{x}(t)$ $=F_{\theta}(t,x(t))$ (S5)
In this work, we resort to the Diffrax library [25, 67] for numerical integration in Jax to solve the above ODE. This powerful library provides efficient implementations of many fixed and adaptive timestep solvers [68, 69]. It also allows easily switching between different modes for backpropagation: Discretize-then-optimize (also referred to as unrolling) and optimize-then-discretize (also referred to as backsolving) [70, 25]. Overall, the choice of solver, stepsize controller and differentiation technique is typically application-dependent. Hence, we implement the general Diffrax integrator as an easy-to-use alternative for the existing fixed-stepsize integrators in MJX, offering the full flexibility of the Diffrax library to the user. Notably, we adjust the Diffrax solvers to accommodate for exact integration of quaternions and stateful actuators, similar to the Runge-Kutta implementation of MJX, which can help to reduce the number of required integration steps.
Figure S2: DiffMJX and CFD add tuning knobs to MJX’s gradient computation. DiffMJX adds support for adaptive integration atop MJX. The integrator’s error tolerances allows to trade computational speed for improved gradient accuracy. The number of checkpointing steps trades computation speed for improved GPU memory consumption. CFD adds parameters that set the distance and magnitude at which artificial contact forces are applied in the gradient computation.
Discretize-then-optimize.
The first notable differentiation mode is discretize-then-optimize. It is the result of discretizing the forward ODE with a numerical integrator and then computing the gradients of the discretized forward ODE by unrolling the computation graph. This approach aligns with the paradigm of differentiable programming used in autodifferentiation libraries such as JAX, hence this is also the approach taken in standard MJX. The problem here is that the memory requirements scale linearly with the number of solver steps. For many applications this is a serious bottleneck in standard MJX, and the problem is exacerbated when relying on adaptive integration, as it typically involves storing even more substeps. Fortunately, Diffrax already implements an optimal gradient checkpointing scheme that allows reducing memory requirements from $O(n)$ to $O(1)$ in exchange for increasing runtime from $O(n)$ to $O(n\log n)$, where $n$ is the maximum number of solver steps.
Another potential issue is that discretize-then-optimize inherently computes gradients of the discretized dynamics, rather than the true continuous dynamics. This can lead to interesting failure cases in which even in the limit of the stepsize going to zero, the computed gradient is different from the true gradient of the continuous dynamics. This is the result of the basic fact that the convergence of a parameterized function (the discretized ODE) to a limit (the continuous ODE) does not imply the convergence of its gradient. Note that this is not a mere mathematical corner case, instead it is the underlying reason for time-of-impact oscillations as described in Figure 5 and [7, 8].
Optimize-then-discretize.
Fixing this requires (approximately) computing the gradients of the true continuous forward ODE. This is achieved by analytically computing the gradient as the solution to the continuous adjoint equations. Assuming a loss $L=L(x(T))$ only on the final state w.l.o.g., the continuous adjoint equations [25, Theorem 5.2] can be written as:
$a_{x}(T)$ $=\frac{dL}{dx(T)}$ $\dot{a}_{x}(t)$ $=-a_{x}(t)^{\top}\frac{\partial F_{\theta}}{\partial x}(t,x(t))$ (S6)
$a_{\theta}(T)$ $=0$ $\dot{a}_{\theta}(t)$ $=-a_{x}(t)^{\top}\frac{\partial F_{\theta}}{d\theta}(t,x(t))$ (S7)
B.3 Contacts from distance with the straight-through-trick
To implement contacts-from-distance only on the backward pass, we use the straight-through trick on the ODE level:
$\dot{x}(t)$ $=\operatorname*{sg}(F_{\theta}(t,x(t)))+\tilde{F}_{\theta}(t,x(t))-\operatorname*{sg}(\tilde{F}_{\theta}(t,x(t))),$ (S8)
where $\operatorname*{sg}$ is the stop-gradient operator. Here $\tilde{F}$ would be obtained by modifying the reference acceleration and impedance as described in Figure 8. In code this reads as:

### v2

@@ 3.1 Adaptive stepsize integration: Numerical precision on demand
A standard method for integrating ODEs with variable stiffness is adaptive integration. The idea behind adaptive integration is elegant: Two numerical integrators of different orders compute the next state. Their difference provides an estimate of the error. If the error is smaller than a given threshold, the step is accepted; otherwise, the step is rejected and the procedure is repeated with a different stepsize chosen by a feedback controller. For further details on the rich history of adaptive stepsize integration, see e.g. Hairer et al. (2008); Hairer & Wanner (2002); Söderlind (2002; 2003).
We use Diffrax (Kidger, 2021) for efficient numerical integration in JAX, taking advantage of its solver flexibility and multiple backpropagation modes. Notably, as detailed in Section˜C.2, we devoted substantial effort to seamlessly integrating quaternions and stateful actuators, which enables seamless compatibility between Diffrax and MJX while ensuring efficient adaptive integration and backwards compatibility with other MuJoCo libraries.
[…]
Testing our implementation on the cube bounce toy example, we observe in Fig.˜5 that adaptive integration reduces the error in the loss and gradient by multiple orders of magnitude given the same computational budget; more detailed analysis is in Fig.˜17 in the Appendix. Importantly, the adaptive stepsize selection in Diffrax is independent across parallelized environments. Therefore, when vmapping across environments, contacts in one simulation do not slow down other simulations. For experimental verification, see Fig.˜20 in Appendix˜D.
@@
Using an adaptive integrator with MJX eliminates oscillations in the bounce example. However, gradients for some object primitives (capsule, cylinder, box) experience gradient artifacts due to non-differentiable operations in the collision detector arising from discrete case distinctions. We smoothed them with standard proxies, finally leading to the results in the bottom row of Fig.˜4, where analytical gradients nearly match central differences. Henceforth, we refer to MJX with the Diffrax integrator and smoothed collision detection as DiffMJX.
@@
Figure 8: Billiard simulation. Top: Force $F$ acts on the white ball affecting the loss. Bottom: Despite the loss derivative being zero if the balls do not collide, DiffMJX with CFD provides informative non-zero gradients.
4 Contacts from distance with straight-through estimation
While adaptive integration improves gradient accuracy, we now shift attention to computing informative gradients between objects that are not in contact. To illustrate why the computation of such gradients is of fundamental importance for robot learning consider the following example.
Example 0 – Billiard shot:
A billiard table is set up as shown in Fig.˜8 (top). At the first timestep, a force $F$ is exerted on the white ball such that it may hit the black ball. The optimization objective is the distance $L$ between the black ball and the target position. If the balls collide, MJX with adaptive integration yields non-zero gradients $\nabla_{F}L$. However, if $F$ does not cause the balls to touch, then $\nabla_{F}L$ is zero and therefore uninformative for optimization.
In the following, we propose contacts from distance (CFD), a method for computing contact forces for positive signed distances $r$ in the gradient computation of a penalty-based simulation that yields informative gradients even if objects are not in contact. This is accomplished by (i) applying artificial contact forces between non-colliding objects, and (ii) using artificial forces only in the gradient computation to maintain simulation realism.
Creating artificial contact forces.
How to generate artificial contact forces in a penalty-based simulator depends on the respective contact model. Below, as a concrete example, we propose a method tailored to MuJoCo. As discussed in Section˜2, the magnitude of contact forces is determined by the impedance $d(r)$ and position-level reference acceleration $h(r)$. To enable the solver to apply CFD, $d(r)$ is augmented as shown in Fig.˜8. Here, $d(r)$ remains unaltered for $r<0$ and is extended by an additional spline for $r>0$. This continuation is parametrized by solimp-CFD parameters $(d_{c},d_{0},w_{c},m_{c},p_{c})$. By default, the curve smoothly continues MuJoCo’s impedance at $d_{0}$ and tapers off to $d_{c}=0$ to ensure smooth differentiability. The CFD-width $w_{c}$ specifies the distance for which artificial contact forces are generated. Moreover, we soften the reference acceleration $h(r)$ by replacing the ReLU function on the signed distance with a softplus (Fig.˜8). This yields modified contact forces $f_{\text{CFD}}$.
@@ 4 Contacts from distance with straight-through estimation
Naively adding CFD to a simulation produces nonphysical behaviors. As shown in Fig.˜8 (top), the artificial contact forces would cause a quadruped to hover above the ground as if a soft foam mat of thickness $w_{c}$ had been placed on the surface. As significantly altering simulation realism is not an option, we are faced with the question: Can CFD be used to obtain informative contact gradients without affecting simulation realism?
To positively answer this question, we resort to the straight-through-trick on the ODE level:
$\dot{x}(t)$ $=\operatorname*{sg}(F_{\theta}(t,x(t)))+\tilde{F}_{\theta}(t,x(t))-\operatorname*{sg}(\tilde{F}_{\theta}(t,x(t))),$ (4)
where $\operatorname*{sg}$ is the stop-gradient operator of an automatic differentiation library. Here $F$ denotes the original ODE obtained from MJX or DiffMJX and $\tilde{F}$ denotes the ODE using CFD via $\dot{v}_{\text{CFD}}=M^{-1}\left(\tau-c+J^{\top}f_{\text{CFD}}\right)$. In JAX this reads as: